Use Mu ̈ller’s method to determine the real and complex roots of
f (x) = x^3− x^2 + 2x − 2
1
Expert's answer
2020-11-05T13:53:13-0500

Let "f(x)=x^3-x^2+2x-2"

Use initial guesses "x_0=-2, x_1=0, x_2=2"

"f(x_0)=f(-2)=-18"

"f(x_1)=f(0)=-2"

"f(x_2)=f(2)=6"

"h_0=x_1-x_0=0-(-2)=2"

"h_1=x_2-x_1=2-0=2"

"\delta_0=\dfrac{f(x_1)-f(x_0)}{h_0}=\dfrac{-2-(-18)}{2}=8" "\delta_1=\dfrac{f(x_2)-f(x_1)}{h_1}=\dfrac{6-(-2)}{2}=4"

"a=\dfrac{\delta_1-\delta_0}{h_1+h_0}=\dfrac{4-8}{2+2}=-1"

"b=a\times h_1+\delta_1=-1\times2+4=2"

"c=f(x_2)=f(2)=6"


"x_3=x_2+\dfrac{-2c}{b\pm\sqrt{b^2-4ac}}"

"x_3=2+\dfrac{-2(6)}{2+\sqrt{2^2-4(-1)(6)}}=0.35424869"

Relative percent error


"\varepsilon_{a^1}=|\dfrac{x_3-x_2}{x_3}|\cdot100\%"

"\begin{matrix}
 n & 1 & 2 & 3 \\
 x_0 & -2 & 0 & 2 \\
 x_1 & 0 & 2 & 0.354249 \\
 x_2 & 2 & 0.354249 & 0.828618 \\
 f(x_0) & -18 & -2 & 6 & \\
 f(x_1) & -2 & 6 & -1.3725396 \\
 f(x_2) & 6 & -1.3725396 & -0.460436 \\
 a & -1 & 1.354249 & 2.182867 \\
 b & 2 & 2.250984 & 2.958255 \\
 c & 6 & -1.3725396 & -0.460436 \\
 x_3 & 0.354249 & 0.828618 & 0.969597 \\
 \varepsilon_{a^n} & 464.575131 & 57.248251 & 14.539952
\end{matrix}"


"\begin{matrix}
 n & 4 & 5 & 6 \\
 x_0 & 0.354249 & 0.828618 & 0.969597 \\
 x_1 & 0.828618 & 0.969597 & 1.001174 \\
 x_2 & 0.969597 & 1.001174 & 0.999998 \\
 f(x_0) & -1.372539 & -0.460436 & -0.089389 \\
 f(x_1) & -0.460436 & -0.089389 & 0.003525 \\
 f(x_2) & -0.089389 & 0.003525 & -0.000006 \\
 a & 1.152467 & 1.799389 & 1.970769\\
 b & 2.794410 & 2.999252 & 3.000028 \\
 c & -0.089389 & 0.003525 & -0.000006 \\
 x_3 & 1.001174 & 0.999998 & 1 \\
 \varepsilon_{a^n} & 3.154012 & 0.117612 & 0.000204
\end{matrix}"

Approximate root of the equation "x^3-x^2+2x-2=0" using Muller method is "1" (After 6 iterations)